# Ultrahigh-energy $\gamma$-ray emission associated with the tail of a bow-shock pulsar wind nebula

### v1

@@ Abstract
Zhen Cao1,2,3, F. Aharonian4,5, Axikegu6, Y.X. Bai1,3, Y.W. Bao7, D. Bastieri8, X.J. Bi1,2,3, Y.J. Bi1,3, W. Bian9, A.V. Bukevich10, Q. Cao11, W.Y. Cao12, Zhe Cao13,12, J. Chang14, J.F. Chang1,3,13, A.M. Chen9, E.S. Chen1,2,3, H.X. Chen15, Liang Chen16, Lin Chen6, Long Chen6, M.J. Chen1,3, M.L. Chen1,3,13, Q.H. Chen6, S. Chen17, S.H. Chen1,2,3, S.Z. Chen1,3, T.L. Chen18, Y. Chen7, N. Cheng1,3, Y.D. Cheng1,2,3, M.C. Chu19, M.Y. Cui14, S.W. Cui11, X.H. Cui20, Y.D. Cui21, B.Z. Dai17, H.L. Dai1,3,13, Z.G. Dai12, Danzengluobu18, X.Q. Dong1,2,3, K.K. Duan14, J.H. Fan8, Y.Z. Fan14, J. Fang17, J.H. Fang15, K. Fang1,3, C.F. Feng22, H. Feng1, L. Feng14, S.H. Feng1,3, X.T. Feng22, Y. Feng15, Y.L. Feng18, S. Gabici23, B. Gao1,3, C.D. Gao22, Q. Gao18, W. Gao1,3, W.K. Gao1,2,3, M.M. Ge17, T.T. Ge21, L.S. Geng1,3, G. Giacinti9, G.H. Gong24, Q.B. Gou1,3, M.H. Gu1,3,13, F.L. Guo16, J. Guo24, X.L. Guo6, Y.Q. Guo1,3, Y.Y. Guo14, Y.A. Han25, O.A. Hannuksela19, M. Hasan1,2,3, H.H. He1,2,3, H.N. He14, J.Y. He14, Y. He6, Y.K. Hor21, B.W. Hou1,2,3, C. Hou1,3, X. Hou26, H.B. Hu1,2,3, Q. Hu12,14, S.C. Hu1,3,27, C. Huang7, D.H. Huang6, T.Q. Huang1,3, W.J. Huang21, X.T. Huang22, X.Y. Huang14, Y. Huang1,2,3, Y.Y. Huang7, X.L. Ji1,3,13, H.Y. Jia6, K. Jia22, H.B. Jiang1,3, K. Jiang13,12, X.W. Jiang1,3, Z.J. Jiang17, M. Jin6, M.M. Kang28, I. Karpikov10, D. Khangulyan1,3, D. Kuleshov10, K. Kurinov10, B.B. Li11, C.M. Li7, Cheng Li13,12, Cong Li1,3, D. Li1,2,3, F. Li1,3,13, H.B. Li1,3, H.C. Li1,3, Jian Li12, Jie Li1,3,13, K. Li1,3, S.D. Li16,2, W.L. Li22, W.L. Li9, X.R. Li1,3, Xin Li13,12, Y.Z. Li1,2,3, Zhe Li1,3, Zhuo Li29, E.W. Liang30, Y.F. Liang30, S.J. Lin21, B. Liu12, C. Liu1,3, D. Liu22, D.B. Liu9, H. Liu6, H.D. Liu25, J. Liu1,3, J.L. Liu1,3, M.Y. Liu18, R.Y. Liu7, S.M. Liu6, W. Liu1,3, Y. Liu8, Y.N. Liu24, Q. Luo21, Y. Luo9, H.K. Lv1,3, B.Q. Ma29, L.L. Ma1,3, X.H. Ma1,3, J.R. Mao26, Z. Min1,3, W. Mitthumsiri31, H.J. Mu25, Y.C. Nan1,3, A. Neronov23, K.C.Y. Ng19, L.J. Ou8, P. Pattarakijwanich31, Z.Y. Pei8, J.C. Qi1,2,3, M.Y. Qi1,3, B.Q. Qiao1,3, J.J. Qin12, A. Raza1,2,3, D. Ruffolo31, A. Sáiz31, M. Saeed1,2,3, D. Semikoz23, L. Shao11, O. Shchegolev10,32, X.D. Sheng1,3, F.W. Shu33, H.C. Song29, Yu.V. Stenkin10,32, V. Stepanov10, Y. Su14, D.X. Sun12,14, Q.N. Sun6, X.N. Sun30, Z.B. Sun34, J. Takata35, P.H.T. Tam21, Q.W. Tang33, R. Tang9, Z.B. Tang13,12, W.W. Tian2,20, L.H. Wan21, C. Wang34, C.B. Wang6, G.W. Wang12, H.G. Wang8, H.H. Wang21, J.C. Wang26, Kai Wang7, Kai Wang35, L.P. Wang1,2,3, L.Y. Wang1,3, P.H. Wang6, R. Wang22, W. Wang21, X.G. Wang30, X.Y. Wang7, Y. Wang6, Y.D. Wang1,3, Y.J. Wang1,3, Z.H. Wang28, Z.X. Wang17, Zhen Wang9, Zheng Wang1,3,13, D.M. Wei14, J.J. Wei14, Y.J. Wei1,2,3, T. Wen17, C.Y. Wu1,3, H.R. Wu1,3, Q.W. Wu35, S. Wu1,3, X.F. Wu14, Y.S. Wu12, S.Q. Xi1,3, J. Xia12,14, G.M. Xiang16,2, D.X. Xiao11, G. Xiao1,3, Y.L. Xin6, Y. Xing16, D.R. Xiong26, Z. Xiong1,2,3, D.L. Xu9, R.F. Xu1,2,3, R.X. Xu29, W.L. Xu28, L. Xue22, D.H. Yan17, J.Z. Yan14, T. Yan1,3, C.W. Yang28, C.Y. Yang26, F. Yang11, F.F. Yang1,3,13, L.L. Yang21, M.J. Yang1,3, R.Z. Yang12, W.X. Yang8, Y.H. Yao1,3, Z.G. Yao1,3, L.Q. Yin1,3, N. Yin22, X.H. You1,3, Z.Y. You1,3, Y.H. Yu12, Q. Yuan14, H. Yue1,2,3, H.D. Zeng14, T.X. Zeng1,3,13, W. Zeng17, M. Zha1,3, B.B. Zhang7, F. Zhang6, H. Zhang9, H.M. Zhang7, H.Y. Zhang17, J.L. Zhang20, Li Zhang17, P.F. Zhang17, P.P. Zhang12,14, R. Zhang14, S.B. Zhang2,20, S.R. Zhang11, S.S. Zhang1,3, X. Zhang7, X.P. Zhang1,3, Y.F. Zhang6, Yi Zhang1,14, Yong Zhang1,3, B. Zhao6, J. Zhao1,3, L. Zhao13,12, L.Z. Zhao11, S.P. Zhao14, X.H. Zhao26, F. Zheng34, W.J. Zhong7, B. Zhou1,3, H. Zhou9, J.N. Zhou16, M. Zhou33, P. Zhou7, R. Zhou28, X.X. Zhou1,2,3, X.X. Zhou6, B.Y. Zhu12,14, C.G. Zhu22, F.R. Zhu6, H. Zhu20, K.J. Zhu1,2,3,13, Y.C. Zou35, X. Zuo1,3,
(The LHAASO Collaboration)
@@ 3 TIANFU Cosmic Ray Research Center, Chengdu, Sichuan, China
4 Dublin Institute for Advanced Studies, 31 Fitzwilliam Place, 2 Dublin, Ireland
@@ 10 Institute for Nuclear Research of Russian Academy of Sciences, 117312 Moscow, Russia
11 Hebei Normal University, 050024 Shijiazhuang, Hebei, China
@@ 15 Research Center for Astronomical Computing, Zhejiang Laboratory, 311121 Hangzhou, Zhejiang, China
16 Key Laboratory for Research in Galaxies and Cosmology, Shanghai Astronomical Observatory, Chinese Academy of Sciences, 200030 Shanghai, China
@@ 23 APC, Université Paris Cité, CNRS/IN2P3, CEA/IRFU, Observatoire de Paris, 119 75205 Paris, France
24 Department of Engineering Physics & Department of Astronomy, Tsinghua University, 100084 Beijing, China
@@ 28 College of Physics, Sichuan University, 610065 Chengdu, Sichuan, China
29 School of Physics, Peking University, 100871 Beijing, China

### v2

@@ Abstract
Zhen Cao1,2,3, F. Aharonian3,4,5,6, Y.X. Bai1,3, Y.W. Bao7, D. Bastieri8, X.J. Bi1,2,3, Y.J. Bi1,3, W. Bian7, A.V. Bukevich9, C.M. Cai10, W.Y. Cao4, Zhe Cao11,4, J. Chang12, J.F. Chang1,3,11, A.M. Chen7, E.S. Chen1,3, H.X. Chen13, Liang Chen14, Long Chen10, M.J. Chen1,3, M.L. Chen1,3,11, Q.H. Chen10, S. Chen15, S.H. Chen1,2,3, S.Z. Chen1,3, T.L. Chen16, X.B. Chen17, X.J. Chen10, Y. Chen17, N. Cheng1,3, Y.D. Cheng1,2,3, M.C. Chu18, M.Y. Cui12, S.W. Cui19, X.H. Cui20, Y.D. Cui21, B.Z. Dai15, H.L. Dai1,3,11, Z.G. Dai4, Danzengluobu16, Y.X. Diao10, X.Q. Dong1,2,3, K.K. Duan12, J.H. Fan8, Y.Z. Fan12, J. Fang15, J.H. Fang13, K. Fang1,3, C.F. Feng22, H. Feng1, L. Feng12, S.H. Feng1,3, X.T. Feng22, Y. Feng13, Y.L. Feng16, S. Gabici23, B. Gao1,3, C.D. Gao22, Q. Gao16, W. Gao1,3, W.K. Gao1,2,3, M.M. Ge15, T.T. Ge21, L.S. Geng1,3, G. Giacinti7, G.H. Gong24, Q.B. Gou1,3, M.H. Gu1,3,11, F.L. Guo14, J. Guo24, X.L. Guo10, Y.Q. Guo1,3, Y.Y. Guo12, Y.A. Han25, O.A. Hannuksela18, M. Hasan1,2,3, H.H. He1,2,3, H.N. He12, J.Y. He12, X.Y. He12, Y. He10, S. Hernández-Cadena7, Y.K. Hor21, B.W. Hou1,2,3, C. Hou1,3, X. Hou26, H.B. Hu1,2,3, S.C. Hu1,3,27, C. Huang17, D.H. Huang10, J.J. Huang1,2,3, T.Q. Huang1,3, W.J. Huang21, X.T. Huang22, X.Y. Huang12, Y. Huang1,3,27, Y.Y. Huang17, X.L. Ji1,3,11, H.Y. Jia10, K. Jia22, H.B. Jiang1,3, K. Jiang11,4, X.W. Jiang1,3, Z.J. Jiang15, M. Jin10, S. Kaci7, M.M. Kang28, I. Karpikov9, D. Khangulyan1,3, D. Kuleshov9, K. Kurinov9, B.B. Li19, Cheng Li11,4, Cong Li1,3, D. Li1,2,3, F. Li1,3,11, H.B. Li1,2,3, H.C. Li1,3, Jian Li4, Jie Li1,3,11, K. Li1,3, L. Li29, R.L. Li12, S.D. Li14,2, T.Y. Li7, W.L. Li7, X.R. Li1,3, Xin Li11,4, Y.Z. Li1,2,3, Zhe Li1,3, Zhuo Li30, E.W. Liang31, Y.F. Liang31, S.J. Lin21, B. Liu4, C. Liu1,3, D. Liu22, D.B. Liu7, H. Liu10, H.D. Liu25, J. Liu1,3, J.L. Liu1,3, J.R. Liu10, M.Y. Liu16, R.Y. Liu17, S.M. Liu10, W. Liu1,3, X. Liu10, Y. Liu8, Y. Liu10, Y.N. Liu24, Y.Q. Lou24, Q. Luo21, Y. Luo7, H.K. Lv1,3, B.Q. Ma25,30, L.L. Ma1,3, X.H. Ma1,3, J.R. Mao26, Z. Min1,3, W. Mitthumsiri32, G.B. Mou33, H.J. Mu25, Y.C. Nan1,3, A. Neronov23, K.C.Y. Ng18, M.Y. Ni12, L. Nie10, L.J. Ou8, P. Pattarakijwanich32, Z.Y. Pei8, J.C. Qi1,2,3, M.Y. Qi1,3, J.J. Qin4, A. Raza1,2,3, C.Y. Ren12, D. Ruffolo32, A. Sáiz32, M. Saeed1,2,3, D. Semikoz23, L. Shao19, O. Shchegolev9,34, Y.Z. Shen17, X.D. Sheng1,3, Z.D. Shi4, F.W. Shu29, H.C. Song30, Yu.V. Stenkin9,34, V. Stepanov9, Y. Su12, D.X. Sun4,12, H. Sun22, Q.N. Sun1,3, X.N. Sun31, Z.B. Sun35, N.H. Tabasam22, J. Takata36, P.H.T. Tam21, H.B. Tan17, Q.W. Tang29, R. Tang7, Z.B. Tang11,4, W.W. Tian2,20, C.N. Tong17, L.H. Wan21, C. Wang35, G.W. Wang4, H.G. Wang8, H.H. Wang21, J.C. Wang26, K. Wang30, Kai Wang17, Kai Wang36, L.P. Wang1,2,3, L.Y. Wang1,3, L.Y. Wang19, R. Wang22, W. Wang21, X.G. Wang31, X.J. Wang10, X.Y. Wang17, Y. Wang10, Y.D. Wang1,3, Z.H. Wang28, Z.X. Wang15, Zheng Wang1,3,11, D.M. Wei12, J.J. Wei12, Y.J. Wei1,2,3, T. Wen15, S.S. Weng33, C.Y. Wu1,3, H.R. Wu1,3, Q.W. Wu36, S. Wu1,3, X.F. Wu12, Y.S. Wu4, S.Q. Xi1,3, J. Xia4,12, J.J. Xia10, G.M. Xiang14,2, D.X. Xiao19, G. Xiao1,3, Y.L. Xin10, Y. Xing14, D.R. Xiong26, Z. Xiong1,2,3, D.L. Xu7, R.F. Xu1,2,3, R.X. Xu30, W.L. Xu28, L. Xue22, D.H. Yan15, J.Z. Yan12, T. Yan1,3, C.W. Yang28, C.Y. Yang26, F.F. Yang1,3,11, L.L. Yang21, M.J. Yang1,3, R.Z. Yang4, W.X. Yang8, Y.H. Yao1,3, Z.G. Yao1,3, X.A. Ye12, L.Q. Yin1,3, N. Yin22, X.H. You1,3, Z.Y. You1,3, Y.H. Yu4, Q. Yuan12, H. Yue1,2,3, H.D. Zeng12, T.X. Zeng1,3,11, W. Zeng15, M. Zha1,3, B.B. Zhang17, B.T. Zhang1,3, F. Zhang10, H. Zhang7, H.M. Zhang31, H.Y. Zhang15, J.L. Zhang20, Li Zhang15, P.F. Zhang15, P.P. Zhang4,12, R. Zhang12, S.R. Zhang19, S.S. Zhang1,3, W.Y. Zhang19, X. Zhang33, X.P. Zhang1,3, Yi Zhang1,12, Yong Zhang1,3, Z.P. Zhang4, J. Zhao1,3, L. Zhao11,4, L.Z. Zhao19, S.P. Zhao12, X.H. Zhao26, Z.H. Zhao4, F. Zheng35, W.J. Zhong17, B. Zhou1,3, H. Zhou7, J.N. Zhou14, M. Zhou29, P. Zhou17, R. Zhou28, X.X. Zhou1,2,3, X.X. Zhou10, B.Y. Zhu4,12, C.G. Zhu22, F.R. Zhu10, H. Zhu20, K.J. Zhu1,2,3,11, Y.C. Zou36, X. Zuo1,3, (The LHAASO Collaboration)
@@ 4 University of Science and Technology of China, 230026 Hefei, Anhui, China
5 Yerevan State University, 1 Alek Manukyan Street, Yerevan 0025, Armeni a
@@ 13 Research Center for Astronomical Computing, Zhejiang Laboratory, 311121 Hangzhou, Zhejiang, China
14 Shanghai Astronomical Observatory, Chinese Academy of Sciences, 200030 Shanghai, China
@@ 18 Department of Physics, The Chinese University of Hong Kong, Shatin, New Territories, Hong Kong, China
19 Hebei Normal University, 050024 Shijiazhuang, Hebei, China
@@ 23 APC, Université Paris Cité, CNRS/IN2P3, CEA/IRFU, Observatoire de Paris, 119 75205 Paris, France
24 Department of Engineering Physics & Department of Physics & Department of Astronomy, Tsinghua University, 100084 Beijing, China
@@ 28 College of Physics, Sichuan University, 610065 Chengdu, Sichuan, China
30 School of Physics & Kavli Institute for Astronomy and Astrophysics, Peking University, 100871 Beijing, China
@@ 32 Department of Physics, Faculty of Science, Mahidol University, Bangkok 10400, Thailand
33 School of Physics and Technology, Nanjing Normal University, 210023 Nanjing, Jiangsu, China